# Detecting Multimedia Generated by Large AI Models: A Survey

### v1

@@ 3.2 Image
Figure 6: Illustrations of pure detection methodologies for LAIM-generated image.
@@ 3.2.1 Pure Detection
$\clubsuit$ Physical/Physiological-based Methods. Physical-based methods detect DM-generated images by examining inconsistencies with real-world physics, such as lighting and reflections. Physiologically-based methods, on the other hand, investigate the semantic aspects of human faces [222], including cues such as symmetry, iris color, pupil shapes, skin, etc., see Fig. 6 a). These methods have stronger interpretability than data-driven methods, which have been widely adapted to detect GAN-generated images [223, 224].
@@
$\clubsuit$ Diffuser Fingerprints-based Methods. Every generation mechanism leaves its own unique fingerprints, which can be explored for detection, as shown in Fig. 6 b). In [95], the authors elucidated that CNNs inherently manifest certain image artifacts. These artifacts are harnessed in their Deep Image Fingerprint (DIF) methodology to extract distinct fingerprints from images generated by these networks (GANs and DMs). This approach facilitates the identification of images originating from a specific model or its fine-tuned variants. Complementarily, the studies by Wang et al. [84] and Ma et al. [176] delve into the realm of diffusion models. They have laid the groundwork for detecting DM-generated images by leveraging the deterministic reverse and denoising processes inherent diffusion models.
@@
Though Wang et al. [84] indeed leverages some deterministic aspects, their approach (i.e., DIRE) primarily targets the reconstruction at the initial time step $\mathbf{x}_{0}$. This method potentially overlooks the rich information present in the intermediate stages of the diffusion and reverse diffusion processes. Ma et al. [176] exploits these intermediate steps later. They design Stepwise Error for Diffusion-generated Image Detection (SeDID), particularly focusing on the errors between reverse and denoise samples at specific time steps during the generation process.
@@
$\clubsuit$ Spatial-based Methods. This research collection focuses on mining spatial characteristics and features within images to detect DM-generated content. Each study utilizes different spatial aspects of images, such as texture, gradients, and local intrinsic dimensionality, for detection, see Fig. 6 c). Motivated by the principle that pixels in rich texture regions exhibit more significant fluctuations than those in poor texture regions. Consequently, synthesizing realistic rich texture regions proves to be more challenging for existing generative models. Zhang et al. [94] leverage such texture features and exploit the contrast in inter-pixel correlation between rich and poor texture regions within an image for DM-generated image forensics. Nguyen et al. [174] use gradient-based features to differentiate DM-generated and human-made artwork by Lorenz et al. [175] propose using the lightweight multi Local Intrinsic Dimensionality (multiLID) for effective detection. This approach stems from the principle that neural networks heavily rely on low-dimensional textures [226] and natural images can be represented as mixtures of textures residing on a low-dimensional manifold [227]. Leveraging this concept, multiLID scores are calculated on the lower-dimensional feature-map representations derived from ResNet18. Then a classifier (random forest) is trained on these multiLID scores to distinguish between synthetic and real images. Their proposed multiLID approach exhibits superiority in diffusion detection. However, this solution does not offer good transferability.
@@
$\clubsuit$ Frequency-based Methods. Frequency-based methods analyze the frequency components of an image to extract information that is not readily apparent in the spatial domain, see Fig. 6 d). The initial study by Wolter et al. [171] highlights the limitations of traditional detection methods that primarily utilize spatial domain convolutional neural networks or Fourier transform techniques. In response, they propose an innovative approach using multi-scale wavelet representation, which incorporates elements from both spatial and frequency domains. However, the study notes limited benefits from higher-order wavelets in guided diffusion data, pinpointing this as a potential area for future research. In a recent study, Xi et al. [173] develop a dual-stream network that combines texture information and low-frequency analysis. This approach identifies artificial images by focusing on low-frequency forgeries and detailed texture information. It has shown to be efficient across various image resolutions and databases. [172] introduces AUSOME to detect DALL-E 2 images by performing a spectral comparison of Discrete Fourier Transform (DFT) and Discrete Cosine Transform (DCT). Bammey [70] uses the cross-difference filter on the image to highlight the frequency artifacts from DM-generated images. It shows some generalization ability, as well as robustness to JPEG compression.
@@ 3.2.2 Beyond Detection
$\clubsuit$ Interpretability.This picture looks like someone I know, and if the AI algorithm tells it is fake or real, then what is the reasoning and should I trust? " Detection with interpretability is working towards solving such question, see Fig. 7 c). In pursuit of creating interpretable detectors, Aghasanli et al. [162] propose a method for DM-generated image detection, taking advantage of features from fine-tuned ViTs combined with existing classifiers such as Support Vector Machines.
@@
Fully-supervised. Localization in a fully-supervised setting requires detectors to be explicitly trained for localization as a segmentation problem with localization mask label. Guo et al. [83] propose a hierarchical fine-grained Image Forgery Detection and Localization (IFDL) framework with three components: a multi-branch feature extractor, localization, and classification modules. The localization module segments pixel-level forgery regions, while the classification module detects image-level forgery. [161] centers on detecting and segmenting artifact areas that are only noticeable to human perception, not full manipulation region. In addition to providing a pixel-level localization map, the study by Guillaro et al. [87] also offers an integrity score to assist users in their analysis.
@@
$\clubsuit$ Robustness. A robust detector is strategically developed to counteract different attacks consisting of intentionally designed perturbations or noise, as shown in Fig. 7 e). It is also designed to maintain consistent detection performance when exposed to real-world conditions, such as dealing with image compression and blurring.
Spatial-based. Ju et al. [157] propose a synthesized image dataset that can contain diverse post-processing techniques to simulate real-world applications. They further introduce the Global and Local Feature Fusion Framework to enhance image detection by integrating multi-scale global features with detailed local features. Another work presented by Xu et al. [158] also leverages multi-level feature representation for enhancing robustness. Different from the above works, [159] highlights the effectiveness of computing local statistics, as opposed to global statistics, in distinguishing digital camera images from DM-generated images. This method produces promising robustness to image resizing and JPEG compression.
Frequency-based. Contrasting with the approaches that mine spatial features, Hooda et al. [156] leverage frequency features through a unique disjoint ensembling approach. They utilize redundant information in the frequency feature space of DM-generated images to generate disjoint ensembles for adversarial DM-generated image detection.
@@
Other works [154, 62] empirically study human face detection. [154] focuses on detecting profile photos of online accounts, finding that a detector (EfficientNet-B1 [228]) only trained on face images completely fails when detecting non-face images, and the detector can learn the semantic-level artifacts because it remains highly accurate even when the image is compressed to a low resolution. [62] investigates the difficulty of non-expert humans in recognizing fake faces and further concludes that trained models outperform non-expert human users, which brings out the need for solutions to contrast the spread of disinformation.
@@
Furthermore, [67] centers on the biological images and narrows on detecting synthetic western blot images. They explore a range of detection strategies, such as binary classification and one-class detectors, and notably, they do this without utilizing synthetic images in training. Their findings demonstrate the successful identification of synthetic biological images. [155] is the first work to detect authentic-looking handwriting generated by DMs. Their experiments reveal that the strongest discriminative features come from real-world artifacts in genuine handwriting that are not reproduced by generative methods.
@@ 3.3 Video
Recently, there has been limited research focused on detecting videos generated by LAIMs. This is primarily because video generation techniques are more complex than those for images. Additionally, as outlined in Section 2, video generation technology is still in its nascent stages compared with image generation. To the best of our knowledge, only one work [152] related to this area, see Fig. 8, and we classify it into the Beyond Detection category.
Figure 8: Illustration of detection methodology in generalization task for LAIM-generated video.

### v4

@@ 3.2 Image
Figure 6: Illustrations of pure detection methodologies for LAIM-generated image. In a), The red circles highlight regions that violate physical or physiological principles, indicating artifacts commonly found in LAIM-generated images.
@@ 3.2.1 Pure Detection
$\clubsuit$ Watermarking. Recent watermarking methods enable the detection of LAIM-generated images by identifying embedded watermarks introduced during or after the generation process. see Fig. 6 a).
WaDiff [285] embeds user-specific watermarks directly into the generation process using a conditioned input and custom loss functions, enabling detection and identification even in black-box settings. Stable Signature [286] fine-tunes only the latent decoder to embed binary signatures but is limited to latent diffusion models and requires architectural access. In contrast, Tree-Ring Watermarking [287] perturbs the initial noise in Fourier space without retraining, offering broad model compatibility and robustness to common image transformations.
Other approaches include watermark-embedded adversarial examples [288], which embed visible watermarks via adversarial perturbations to prevent diffusion models from mimicking copyrighted content—focusing more on protection than traceability. Finally, Gaussian Shading [289] achieves performance-lossless watermarking by encoding signals into Gaussian latent codes without altering model parameters, offering strong robustness and seamless deployment across diffusion models.
$\clubsuit$ Physical/Physiological-based Methods. Physical-based methods detect DM-generated images by examining inconsistencies with real-world physics, such as lighting and reflections. Physiologically-based methods, on the other hand, investigate the semantic aspects of human faces [346], including cues such as symmetry, iris color, pupil shapes, skin, etc., see Fig. 6 b). These methods have stronger interpretability than data-driven methods, which have been widely adapted to detect GAN-generated images [347, 348].
@@
$\clubsuit$ Diffuser Fingerprints-based Methods. Every generation mechanism leaves its own unique fingerprints, which can be explored for detection, as shown in Fig. 6 c). In [104], the authors elucidated that CNNs inherently manifest certain image artifacts. These artifacts are harnessed in their Deep Image Fingerprint (DIF) methodology to extract distinct fingerprints from images generated by these networks (GANs and DMs). This approach facilitates the identification of images originating from a specific model or its fine-tuned variants. Complementarily, the studies by Wang et al. [93] and Ma et al. [278] delve into the realm of diffusion models. They have laid the groundwork for detecting DM-generated images by leveraging the deterministic reverse and denoising processes inherent diffusion models.
@@
Though Wang et al. [93] indeed leverages some deterministic aspects, their approach (i.e., DIRE) primarily targets the reconstruction at the initial time step $\mathbf{x}_{0}$. This method potentially overlooks the rich information present in the intermediate stages of the diffusion and reverse diffusion processes. Ma et al. [278] exploits these intermediate steps later. They design Stepwise Error for Diffusion generated Image Detection (SeDID), particularly focusing on the errors between reverse and denoise samples at specific time steps during the generation process. Luo et al. [279] further improve efficiency and effectiveness with LaRE², a Latent REconstruction error-guided feature REfinement method. Instead of full inversion, LaRE² computes reconstruction error via single step denoising in the latent space, making it over 8× faster than multi-step approaches like DIRE [93].
Rajan et al. [280] focus on improving fingerprint-based detection by aligning real and fake images more precisely. Instead of generating fake images through the standard denoising pipeline, they reconstruct real images using only the LDM’s autoencoder, capturing artifacts introduced by the decoder while maintaining alignment in resolution, semantics, and format. This allows the detector to focus specifically on decoder induced fingerprints, avoiding spurious correlations. Brokman et al. [281] present a theoretically grounded zero-shot detection method that leverages the score function of pre-trained diffusion models to quantify biases in generated images. Their method defines geometric criteria, curvature and gradient magnitude, on the implicit probability manifold learned by the model. These quantities, when computed over local perturbations of a given image, serve as indicators of whether the image is fake or real.
@@
$\clubsuit$ Spatial-based Methods. This research collection focuses on mining spatial characteristics and features within images to detect DM-generated content. Each study utilizes different spatial aspects of images, such as texture, gradients, and local intrinsic dimensionality, for detection, see Fig. 6 d). Motivated by the principle that pixels in rich texture regions exhibit more significant fluctuations than those in poor texture regions. Consequently, synthesizing realistic rich texture regions proves to be more challenging for existing generative models. Zhang et al. [103] leverage such texture features and exploit the contrast in inter-pixel correlation between rich and poor texture regions within an image for DM-generated image forensics. Nguyen et al. [276] use gradient-based features to differentiate DM-generated and human-made artwork. Lorenz et al. [277] propose using the lightweight multi Local Intrinsic Dimensionality (multiLID) for effective detection. This approach stems from the principle that neural networks heavily rely on low-dimensional textures [350] and natural images can be represented as mixtures of textures residing on a low-dimensional manifold [351]. Leveraging this concept, multiLID scores are calculated on the lower dimensional feature map representations derived from ResNet18. Then a classifier (random forest) is trained on these multiLID scores to distinguish between synthetic and real images. Their proposed multiLID approach exhibits superiority in diffusion detection. However, this solution does not offer good transferability.
@@
$\clubsuit$ Frequency-based Methods. Frequency-based methods analyze the frequency components of an image to extract information that is not readily apparent in the spatial domain, see Fig. 6 e). The initial study by Wolter et al. [272] highlights the limitations of traditional detection methods that primarily utilize spatial domain convolutional neural networks or Fourier transform techniques. In response, they propose an innovative approach using multi-scale wavelet representation, which incorporates elements from both spatial and frequency domains. However, the study notes limited benefits from higher order wavelets in guided diffusion data, pinpointing this as a potential area for future research. In a recent study, Xi et al. [274] develop a dual stream network that combines texture information and low frequency analysis. This approach identifies artificial images by focusing on low frequency forgeries and detailed texture information. It has shown to be efficient across various image resolutions and databases. [273] introduces AUSOME to detect DALL-E 2 images by performing a spectral comparison of Discrete Fourier Transform (DFT) and Discrete Cosine Transform (DCT). Bammey [79] uses the cross difference filter on the image to highlight the frequency artifacts from DM-generated images. It shows some generalization ability, as well as robustness to JPEG compression. Lanzino et al. [275] propose a real-time deepfake detector using Binary Neural Networks, incorporating FFT and LBP features to capture frequency and texture artifacts with minimal accuracy loss.
$\clubsuit$ Distribution-based Methods. Instead of analyzing images in the spatial or frequency domain directly, distribution-based methods model the intrinsic statistical distribution of real images and detect synthetic images as distributional anomalies, see Fig. 6 f). Cozzolino et al. [271] propose a zero-shot detector (ZED) that leverages a lossless image compression model trained only on real images to predict the probability distribution of each pixel given its context. By computing the gap between the actual negative log-likelihood (NLL) and the expected entropy across multiple resolution scales, ZED achieves strong detection performance without requiring any synthetic training data.
@@ 3.2.2 Beyond Detection
More recently, Baraldi et al. [151] proposed CoDE, a contrastive learning framework that enforces global-local similarities to improve generalization across different generated image types. Yuan et al. [266] introduced HRR, a diffusion-based detection model that enhances generalization through multi-scale retrospection and feature refinement modules, mitigating dataset and generator biases. Yan et al. [112] propose AIDE, a hybrid detector that combines frequency-domain features and CLIP-based semantic features. AIDE achieves state-of-the-art results across both general (GenImage [81], AIGCD [103]) and challenging (Chameleon [112]) benchmarks. Cheng et al. [267] address the catastrophic forgetting problem in incremental face forgery detection, which directly impacts generalization to new forgery types. Their method, SUR-LID, introduces Sparse Uniform Replay (SUR) to approximate the latent feature distribution of past tasks and Latent-space Incremental Detector (LID) to align new features “brick by brick.” The framework enables continual learning across datasets and forgery types while retaining high detection accuracy. To further enhance generalization, Guillaro et al. [268] propose B-Free, a bias-free training paradigm that generates self-conditioned fake images from real ones to avoid semantic and compression biases. Combined with content augmentation, B-Free significantly improves generalization across 27 generative models. Meanwhile, Zheng et al. [269] identify cross-scene generalization as a key challenge, where detectors struggle on images with unseen semantic content due to overfitting to “semantic artifacts.” They address this by introducing a patch-shuffling strategy and a patch-based classifier that disrupts global scene cues.
$\clubsuit$ Interpretability.This picture looks like someone I know, and if the AI algorithm tells it is fake or real, then what is the reasoning and should I trust? " Detection with interpretability is working towards solving such question, see Fig. 7 c). In pursuit of creating interpretable detectors, Aghasanli et al. [257] propose a method for DM-generated image detection, taking advantage of features from fine-tuned ViTs combined with existing classifiers such as Support Vector Machines. This approach is notable for interpreting the model behavior through prototypes by analyzing SVM’s closest support vectors. using SVM’s closest support vectors, enabling a unique form of interpretability-through prototypes. Color distributions offer a rich and interpretable cue for synthetic image detection. Uhlenbrock et al. [258] show that perceptual loss functions in diffusion models emphasize luminance over chrominance, leading to unnatural relationships between color channels. By transforming images into multiple color spaces and computing residual-based co-occurrence statistics, their simple random forest detector achieves strong generalization across generators and post-processing variants.
@@
Fully-supervised. Localization in a fully-supervised setting requires detectors to be explicitly trained for localization as a segmentation problem with localization mask label. Guo et al. [92] propose a hierarchical fine-grained Image Forgery Detection and Localization (IFDL) framework with three components: a multi-branch feature extractor, localization, and classification modules. The localization module segments pixel-level forgery regions, while the classification module detects image-level forgery. [255] centers on detecting and segmenting artifact areas that are only noticeable to human perception, not full manipulation region. In addition to providing a pixel-level localization map, the study by Guillaro et al. [96] also offers an integrity score to assist users in their analysis. Li et al. [256] introduce UnionFormer, a unified-learning transformer framework for simultaneous detection and localization. It incorporates three complementary views: RGB, noise, and object consistency. The proposed BSFI-Net enhances boundary sensitivity and captures artifacts across scales, while a unified transformer integrates multi-view clues to yield both detection results and a pixel-level manipulation mask.
@@
$\clubsuit$ Robustness. A robust detector is strategically developed to counteract different attacks consisting of intentionally designed perturbations or noise, as shown in Fig. 7 e). It is also designed to maintain consistent detection performance, for example, if the input image is fake, a robust detector should consistently predict it as fake, regardless of whether the input has been subjected to attacks such as compression, blurring, or other real-world degradations.
Adversarial Attack Robustness. De Rosa et al.[253] study the adversarial robustness of CLIP-based forensic detectors compared to CNN-based ones. They show that while both architectures are vulnerable to white-box attacks, adversarial perturbations exhibit different frequency-domain patterns between CNNs and CLIP-based models, making attacks less transferable across architectures. Hooda et al.[252] propose D4, an ensemble-based detector partitioning frequency components across models. This disjoint design reduces the adversarial subspace dimensionality and achieves strong robustness against black-box adversarial attacks, outperforming conventional ensemble defenses
Post-Processing Robustness. Ju et al. [249] address robustness against post-processing such as JPEG compression, Gaussian blurring. They propose GLFF, a Global and Local Feature Fusion framework that combines multi-scale global and subtle local features, achieving improved performance on the challenging DF3 dataset with diverse post-processing operations. Xu et al. [250] also leverages multi-level feature representation for enhancing robustness. Different from the above works, [251] introduces a non-deep learning method based on Bayer pattern statistics and local variance features to detect diffusion-generated images. Their method remains effective under image resizing and JPEG compression, showing resilience to perturbations.
@@
Other works [247, 71] empirically study human face detection. [247] focuses on detecting profile photos of online accounts, finding that a detector (EfficientNet-B1 [352]) only trained on face images completely fails when detecting non-face images, and the detector can learn the semantic level artifacts because it remains highly accurate even when the image is compressed to a low resolution. [71] investigates the difficulty of non-expert humans in recognizing fake faces and further concludes that trained models outperform non-expert human users, which brings out the need for solutions to contrast the spread of disinformation.
@@
Furthermore, [76] centers on the biological images and narrows on detecting synthetic western blot images. They explore a range of detection strategies, such as binary classification and one-class detectors, and notably, they do this without utilizing synthetic images in training. Their findings demonstrate the successful identification of synthetic biological images. [248] is the first work to detect authentic looking handwriting generated by DMs. Their experiments reveal that the strongest discriminative features come from real-world artifacts in genuine handwriting that are not reproduced by generative methods.
3.2.3 Analysis
$\clubsuit$ Insights. Across the surveyed detection methods, several shared insights emerge regarding the vulnerabilities of DM-generated images and the design principles of effective detectors. Despite the rapid improvement of diffusion models, their outputs still exhibit subtle artifacts that remain detectable across multiple domains. These include geometric inconsistencies, such as perspective distortions and unnatural reflections [282, 283], as well as low-level generation artifacts left by decoder structures or denoising steps [280, 93]. Many successful detection frameworks exploit the deterministic behavior of diffusion processes, using inversion and reconstruction-based strategies to quantify deviations between real and generated images [278, 281]. In parallel, spatial and frequency domain features continue to serve as strong indicators, with studies showing that texture irregularities [103] and abnormal frequency patterns [274, 272] often co-occur in synthetic images. However, generalization remains a persistent challenge: detectors trained on specific generators or domains often fail when applied to unseen models or image categories, especially when semantic content shifts significantly [269, 102]. Moreover, robustness to post-processing and adversarial perturbations is limited, with even strong detectors showing vulnerability to common transformations like JPEG compression, resizing, or blurring [253, 249].
$\clubsuit$ Recommendations. To address these challenges, several design principles can be recommended. First, hybrid detectors that combine spatial, frequency, and latent cues tend to exhibit improved robustness and generalizability [256, 249]. Second, inversion-based approaches that leverage model-induced consistency, such as reconstruction error or decoder fingerprinting, offer interpretable and transferable signals [278, 280]. Third, promoting open-world generalization requires methods to avoid overfitting to dataset-specific or generator-specific patterns. Strategies such as patch-based classification [269], fingerprint regularization [262], or contrastive learning [151] show promise in this direction. Fourth, interpretability remains crucial for user trust and adoption, particularly in forensic applications; color-space analysis and prototype-based reasoning have emerged as promising paths for interpretable detection [257, 258]. Lastly, to ensure real-world applicability, detectors must be evaluated under diverse image transformations and compression schemes, reflecting practical deployment environments like social media or messaging platforms [102, 251].
$\clubsuit$ Social Media. The detection of LAIM-generated images on social media has emerged as a critical research challenge, unlike controlled lab settings, social media platforms introduce compression, resizing, adversarial editing, and a lack of reliable metadata, causing lab-trained detectors to underperform in real-world environments. ODDN [353] tackles the unpaired data problem by training directly on compressed fake images without requiring pristine counterparts, demonstrating the need for explicitly training detectors for robustness against compression artifacts.
To enable real-world deployment, Cavia et al.[354] propose LaDeDa, a lightweight detector suited for real-time use. While it performs well on lab datasets, tests on the WildRF benchmark, comprising real social media images, reveal notable performance drops. A critical reassessment by Yan et al.[112] further shows that even top detectors often misclassify highly realistic LAIM images. Advancing the field, Huang et al.[110] present SIDA, the first framework for both detection and localization of LAIM images on social media. Leveraging large vision-language models and the newly curated SID-Set dataset, SIDA enables not only accurate detection but also pixel-level localization and interpretable explanations.
Key findings across these studies underline that detecting LAIM-generated images on social media is significantly more difficult than in controlled lab environments. Real-world images often undergo aggressive transformations such as JPEG compression, resizing, filtering, and adversarial post-processing, all of which obscure generative artifacts that detectors typically exploit. Moreover, the dominant sources of AI images on social platforms have shifted from earlier GAN models to advanced diffusion models, which produce much higher quality and contextually realistic outputs. Lab-generated datasets, in contrast, are curated, balanced, and captured under idealized conditions, leading to overestimation of detector performance.
@@ 3.3 Video
Recently, there has been limited research focused on detecting videos generated by LAIMs. This is primarily because video generation techniques are more complex than those for images. Additionally, as outlined in Section 2, video generation technology is still in its nascent stages compared with image generation.
Figure 8: Illustration of detection methodology for LAIM-generated video.
3.3.1 Pure Detection
$\clubsuit$ Spatial&Temporal-based Methods. Ji et al. [243] propose DuB3D, a Dual-Branch 3D Transformer that leverages both spatiotemporal and motion cues for LAIM-generated video detection. DuB3D integrates appearance and optical flow branches through a late fusion mechanism. Extensive experiments show its superior performance. He et al. [244] propose a detection method that jointly models local motion inconsistencies and global appearance variations leveraging a frame prediction error module and a temporal transformer over BEiT [355] features. These are fused via a channel attention module to adaptively highlight defects.
@@ 3.3.2 Beyond Detection
$\clubsuit$ Empirical Study. Vahdati et al. [241] conduct a comprehensive study revealing that detectors trained on synthetic images fail to generalize to AI-generated videos. Their experiments show that this failure is not primarily due to compression artifacts (e.g., H.264), but rather stems from fundamentally different forensic traces left by video generators compared to image generators. To address this gap, they demonstrate that synthetic video traces can be learned directly, enabling reliable detection and generator source attribution even after compression. Furthermore, their study highlights that while zero-shot detection of new video generators remains challenging, few-shot learning can effectively adapt detectors to unseen generators with minimal data.
3.3.3 Analysis.
$\clubsuit$ Insights. Although video generation is still evolving, early detection methods already reveal several shared insights about LAIM-generated video artifacts. Most notably, LAIM-generated videos introduce temporal inconsistencies, such as unnatural motion patterns or frame prediction anomalies, which differ from spatial inconsistencies observed in synthetic images [243, 244]. These motion related artifacts require models to incorporate both spatial and temporal reasoning, often through dual-branch or transformer based architectures. Additionally, empirical studies highlight that forensic traces in videos are fundamentally different from those in images, and detectors trained on image data perform poorly on synthetic videos [241]. Generalization remains a central challenge: video detectors often fail on unseen generators due to overfitting to known artifacts or reliance on overly constrained datasets [242].
$\clubsuit$ Recommendations. To address current limitations, detection systems should exploit temporal dynamics and motion inconsistencies unique to synthetic videos. Dual-branch or transformer based models that integrate spatial and motion cues, enhanced by optical flow and temporal attention, have shown promise [243, 244]. Since detectors trained on images often fail on video [241], training on synthetic video data is crucial to capture video-specific forensic signals. Domain adaptation and few-shot fine-tuning can help generalize across generation methods. A good benchmark should cover diverse sources, compression levels, and durations to reflect real-world conditions. Incorporating frame-level localization and temporal consistency analysis will also improve interpretability and traceability.
$\clubsuit$ Social Media. Efforts such as PDID [130] compile real-world political deepfake incidents, emphasizing that the deployment of generative AI in sociopolitical contexts requires not only detection but also content tracking and contextual analysis. In parallel, benchmarks like WildDeepfake [117] reveal that conventional artifact-based detection methods, effective in controlled settings, often fail against social media deepfakes characterized by compression, noise, and diverse manipulation techniques. Yu et al.[356] propose AMSIM, a model that magnifies subtle spatiotemporal inconsistencies through global and multi-timescale local observations, and introduce adversarial data augmentation to simulate unseen social media degradations, improving robustness to wild conditions. Wu et al.[357] develop ITSNet, an interactive two-stream network that enhances deepfake detection by jointly modeling spatiotemporal inconsistencies across RGB and frequency domains, achieving strong performance even under social media distortions and compression artifacts.